A company’s Balance Sheet (in millions)

Assets                                                             Liabilities & Equity

Current                        $  80              

Net Fixed                    $120                            Bonds ($1000 Par)                  130

                                                                        Preferred stocks ($100 Par)   40

Total                           $200                            Common Stock ($1 par)         30

                                                                        Total                                       $200

The company's bonds have 10 years to mature, pay 10% coupon rate semi-annually and comparable bonds' YTM is 14%.

The company’s applicable tax rate is 40%.

The market price of common stock is $10.50 per share.

The common stock is constantly growing at a rate of 6%. The same growth rate is expected to continue for long time in the future. The most recent dividend on the common stock was $1.15.

The flotation cost for new common stocks is 10%.

The market value of the preferred stock is $45 and it pays quarterly dividend of $1.25.

The flotation cost on issuing new preferred stock is 7%

What is the WACC of the company using the book weights of capital structure (Assuming the company will issue new preferred and common stocks)?

8.38%

10.68%

13.69%

12.41%

15.27%
Answer #1

cost of preferred shares = 1.25*4/(45 * (1- 0.07) = 11.95%

cost of new equity = 1.15*1.06/1.5*(1 - 0.10) + 6% = 18.90%

WACC = (30*18.9% + 130*14*(1 - 0.4)% + 40*11.95%) = 10.68%